OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). m + n = 1000 0.05m + 0.06n = 57

OpenStudy (anonymous):

You can do this problem by the elimination method. Multiply m + n = 1000 by .05. You get -.05 -.06 = -50. Add this to .05m + .06n = 57. You get .01n = 7 n = 700 m + 700 = 1000 m = 300 (300,700)